OpenStudy (anonymous):

Could someone explain to me how 3^2 ≡ 2 (mod 7)?

OpenStudy (amistre64):

1 2 3 4 5 6 7 8 9 ^

OpenStudy (anonymous):

Whats confusing me is I thought 3^2 and 2 had to have the same remainder when divided by 7.

OpenStudy (amistre64):

9/7 = 1R2 2/7 = 0R2

OpenStudy (amistre64):

that works as well

OpenStudy (anonymous):

oh!

OpenStudy (anonymous):

i get it. thx!

OpenStudy (across):

mod is just the remainder after division. 9 divided by 7 gives a remainder of 2.

OpenStudy (anonymous):

ok, one last thing. I'm looking at this congruence that says 6≡ -1 (mod 7). Is that because 6/7 = 1 ?

OpenStudy (anonymous):

7/6= 1 , sry